# The Role of Superlattice Phonons in Charge Localization Across Quantum Dot Arrays

### v1

@@ Abstract
Figure 5: The projection of the potential energy surfaces along the primary bending mode coordinate $Q$ in the [100] WZ nonadiabatic (panel (a)) and adiabatic (panel (b)) limits for different chain lengths. Only the first two neighboring states are shown. The solid lines represent the ground donor potential energy surfaces, and the dashed lines represent the excited donor potential energy surfaces. $Q_{0}$ is the ground state reference geometry, and $Q_{\rm eq}$ is the equilibrium geometry for the donor state. Reorganization energy $\lambda$ is the potential energy difference between $Q_{0}$ and $Q_{\rm eq}$ geometry. The splitting between the two potential energy surfaces is approximately two times the electronic coupling $J$ between the diabatic states.
@@
The increase in the reorganization energy with the length of the QD chain and the slowdown of electron transfer across the QD array in the nonadiabatic regime can be explained by considering the adiabatic potential energy surfaces along the superlattice mode that couples strongly to the excess electron, as shown in Fig. 5 panel (a). As the reorganization energy increases with the length of QD assembly, the energy barrier between the neighboring QDs increases. This results in slower electron transfer in WZ [100] structures, as shown in Fig. 2 panel (a). In the adiabatic regime, on the other hand, the energy barrier does not vary significantly with the number of QD units as shown in Fig. 5 panel (b) and the potential barrier is below the thermal energy at $T=300$ K. This is primarily because the large neck width between the QDs leads to rigid superlattice vibrations with higher phonon frequencies compared to the nonadiabatic limit. Supplementary Figures 5-7 show the effects of removing the inter-dot phonons on the transfer dynamics. We found that, in general, inter-dot phonons lead to faster transfer or dephasing rates when the couplings are not too strong. However, as the reorganization energies increase with the QD chain length, the electron eventually becomes localized. The system will transition to the localized regime unless $J$ is increased simultaneously to keep it in the adiabatic limit, maintaining small barriers between neighboring sites. In this case, the superlattice phonons induce faster dephasing, as shown in Fig. 2 panel (d)-(f). In general, superlattice inter-dot phonons lead to larger energy barriers in the nonadiabatic regime while facilitating dephasing in the adiabatic regime.
@@
In conclusion, we have theoretically examined electron transfer in an array of interconnected colloidal QDs. The electronic structure and electron-phonon interactions were computed using semi-empirical pseudopotential methods and the phonon modes were obtained from a force field. Our results reveal that electron transfer dynamics are profoundly influenced by superlattice phonons in the nonadiabatic regime, particularly when the neck connecting neighboring QDs is narrow. Through a simplified site model, we further illustrate that the impact of coupling to superlattice phonons is connected to the symmetry of the electronic wavefunctions and inter-dot vibrations. Specifically, the lack of centrosymmetry in the wurtzite structure results in strong couplings with the inter-dot bending and symmetric stretching modes. The electron-phonon couplings and reorganization energy increase with the number of QDs in the array, thereby localizing electron transfer. In contrast, adiabatic electron transfer in QD arrays with wide neck widths is less influenced by superlattice phonons, enabling rapid transfer across multiple QD sites. These insights are valuable for designing QD arrays or assemblies that exhibit optimal electron transfer properties.
Acknowledgements.
@@
This work was supported in part by the NSF–BSF International Collaboration in the Division of Materials Research program, NSF grant number DMR-2026741 and BSF grant number 2020618. Methods used in this work were provided by the center on “Traversing the death valley separating short and long times in non-equilibrium quantum dynamical simulations of real materials”, which is funded by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research and Office of Basic Energy Sciences, Scientific Discovery through Advanced Computing (SciDAC) program, under Award No. DE-SC0022088.

### v2

@@ I Results and Discussion
Figure 6: The projection of the potential energy surfaces along the primary bending mode coordinate $Q$ in the [100] WZ nonadiabatic (panel (a)) and adiabatic (panel (b)) limits for different QD chain lengths. Only the first two states on the neighboring sites are shown. The solid lines represent the ground donor potential energy surfaces, and the dashed lines represent the excited donor potential energy surfaces. $Q_{0}$ is the ground state reference configuration, and $Q_{\rm eq}$ is the equilibrium configuration for the donor state. Reorganization energy $\lambda$ is the energy difference between $Q_{0}$ and $Q_{\rm eq}$ configuration. The splitting between the two potential energy surfaces is approximately two times the electronic coupling $J$.
@@
Potential energy surfaces along superlattice modes. The increase in the reorganization energy with the length of the QD chain and the slowdown of electron transfer across the QD array in the nonadiabatic regime can be explained by considering the adiabatic potential energy surfaces along the superlattice mode that couples strongly to the excess electron, as shown in Figure 6 panel (a). As the reorganization energy increases with the length of QD assembly, the energy barrier between the neighboring QDs increases. This results in slower electron transfer in WZ [100] structures, as shown in Figure 2 panel (a). In the adiabatic regime, on the other hand, the energy barrier does not vary significantly with the number of QD units as shown in Figure 6 panel (b) and the potential barrier is below the thermal energy at $T=300$ K. This is primarily because the large neck width between the QDs leads to rigid superlattice vibrations with higher phonon frequencies compared to the nonadiabatic limit. Supplementary Figures 5-7 show the effects of removing the inter-dot phonons on the transfer dynamics in both nonadiabatic and adiabatic regimes. We found that, in general, inter-dot phonons lead to faster transfer or dephasing rates when the couplings are not too strong. However, as the reorganization energies increase with the QD chain length, the electron eventually becomes localized. The system will transition to the localized regime unless $J$ is increased simultaneously to keep it in the adiabatic limit, maintaining small barriers between neighboring sites. In this case, the superlattice phonons induce faster dephasing, as shown in Figure 2 panel (d)-(f). In general, superlattice inter-dot phonons lead to larger energy barriers in the nonadiabatic regime while facilitating dephasing in the adiabatic regime.
II Conclusion
@@
In conclusion, we have theoretically examined electron transfer in an array of interconnected colloidal QDs. The electronic structure and electron-phonon interactions were computed using semi-empirical pseudopotential methods, and the nuclear vibrations were obtained from a force field. Our results reveal that electron transfer dynamics are profoundly influenced by superlattice phonons in the nonadiabatic regime, particularly when the neck connecting neighboring QDs is narrow. Through a simplified site model, we further illustrate that the impact of coupling to superlattice phonons is related to the symmetry of the electronic wavefunctions and inter-dot vibrations.
Specifically, the lack of centrosymmetry in the wurtzite structure results in strong couplings with the inter-dot bending and symmetric stretching modes. The electron-phonon couplings and reorganization energy increase with the number of QDs in the array, thereby localizing electron transfer. In contrast, adiabatic electron transfer in QD arrays with wide neck widths is less influenced by superlattice phonons, enabling rapid transfer across multiple QD sites. While the coupling to superlattice bending modes may be suppressed in periodic 2D or 3D QD superlattices attached to a substrate, we believe that our finite-size 1D model can still elucidate the electron dynamics resulting from local vibrations across several QDs in the assembly, whether it is in 1D, 2D, or 3D.
A few aspects neglected in our current approach, which will be the subject of future studies, include disorder resulting from size inhomogeneity of the QD monomers, the effects of ligands and solvent, as well as strain and composition. For the inhomogeneous disorder, we estimated the localization length using a 1D site chain model and found that the population dynamics may be further affected by Anderson localization with a 5% QD size distribution in the nonadiabatic limit, where the coupling between QD sites is small. In contrast, in the adiabatic regime, the estimated Anderson localization length is an order of magnitude longer than the length scales currently studied, resulting in a negligible contribution. Additionally, the presence of surface ligands and solvent can introduce a secondary bath coupled to the electron. Strain and composition effects, such as off-stoichiometry in the QD monomers, can create defect states that influence the transfer dynamics. Future studies will incorporate an atomistic description, particularly focusing on ligand motion and various defect states, to investigate their effects on electron transfer in the QD chain.
III Methods
Semi-Empirical Pseudopotential Method. The semi-empirical pseudopotential method was employed to calculate the quasi-electron Hamiltonian describing the valence electron of the QD arrays. The local screened pseudopotentials were fitted to a functional form in the momentum space as [49, 50, 51, 42]
$\tilde{v}(q)=a_{0}\left[1+a_{4}\operatorname{Tr}\bm{\epsilon}\right]\frac{q^{2}-a_{1}}{a_{2}\exp\left(a_{3}q^{2}\right)-1}$ (6)
where $q$ is the magnitude of momentum vector, $\bm{\epsilon}$ is the strain tensor, and $a_{0}$ to $a_{4}$ are fitting parameters for Cd, Se, and S, which are tabulated in the Supplementary Table 1. The parameters were fitted simultaneously to the band structures and deformation potentials of bulk CdSe and CdS to capture the electronic and vibronic properties of the nanocrystals. Filter diagonalization technique was then applied to obtain quasiparticle electron states $\psi_{i}(\mathbf{r})$ of the Hamiltonian ${h}_{\rm e}(\mathbf{r})$ above the conduction band edge. The calculations were performed on real-space grids with spacings smaller than 0.7 a.u. to ensure that the eigenenergies converged to within $10^{-3}$ meV and that the sequential calculations of electron-phonon couplings were also converged.
To study electron transfer within the QD array, we constructed a set of states $\phi_{n}(\mathbf{r})$ localized on each QD unit, defining a set of donor and acceptor states. The transformation between the delocalized eigenstates $\psi_{i}(\mathbf{r})$ and the localized states $\phi_{n}(\mathbf{r})$ was achieved using the Förster-Boys localization scheme, which maximizes the self-extension criterion [41]. Eigenstates with $s$- and $p$-like envelope functions were included in the localization procedure.
Electron-Phonon Couplings. The first-order atomistic electron-nuclear coupling $V_{nm}^{\mu k}$ between states $\phi_{n}(\mathbf{r})$ and $\phi_{m}(\mathbf{r})$ was calculated by [42]
$V_{nm}^{\mu k}=\bra{\phi_{n}}\frac{\partial h_{\rm e}}{\partial R_{\mu k}}\ket{\phi_{m}},$ (7)
where $R_{\mu k}$ is the atomic coordinates with $\mu=1\dots N_{\rm atom}$ and $k=x,y,z$. In QD arrays with a narrow neck width, the superlattice vibrations have low frequencies. To avoid mixing inter-dot superlattice modes with global translational and rotational motions in this narrow neck limit, we defined a projected electron-nuclear coupling [52] to eliminate those degrees of freedom (see Supplementary Information for details). The QD array structures are relaxed using the 3-body Stillinger-Weber force field [53]. Under the harmonic approximation, the normal mode frequencies $\omega_{\alpha}$ and coordinates $\mathbf{E}^{\alpha}$ can be calculated by diagonalizing the Hessian matrix at the equilibrium configuration. The transformation between normal mode and atomic coordinate is defined as $Q_{\alpha}=\sum_{\mu k}\sqrt{m_{\mu}}(R_{\mu k}-R_{\mu k,0})E_{\mu k}^{\alpha}$, which leads to the definition of electron-phonon couplings
$V_{nm}^{\alpha}=\sum_{\mu,k}\frac{1}{\sqrt{m_{\mu}}}E_{\mu k,\alpha}V_{nm}^{\mu k}.$ (8)
All the parameters are then plugged into the model Hamiltonian in Eqs. 1 for the dynamical calculation.
acknowledgement
@@
This work was supported in part by the NSF–BSF International Collaboration in the Division of Materials Research program, NSF grant number DMR-2026741 and BSF grant number 2020618. Methods used in this work were provided by the center on “Traversing the death valley separating short and long times in non-equilibrium quantum dynamical simulations of real materials”, which is funded by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research and Office of Basic Energy Sciences, Scientific Discovery through Advanced Computing (SciDAC) program, under Award No. DE-SC0022088. E. Rabani acknowledges the Alexander von Humboldt Foundation for the generous Humboldt Research Award.